OpenStudy (anonymous):

How to solve this to find out if its an identity or no solution? 8(h-1)=6h+4+2h

jimthompson5910 (jim_thompson5910):

8(h-1)=6h+4+2h 8h-8=6h+4+2h ... Distribute 8h-8=8h+4 ... Combine like terms. What's next?

OpenStudy (anonymous):

Do you divide 8 by 8 and then also divide 4 by 8?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you subtract 8h from both sides

OpenStudy (anonymous):

oh, and then it leaves you with 8 and 4, so that is no solution right?

jimthompson5910 (jim_thompson5910):

yes since -8 = 4 is a false statement, it's a contradiction so there are no solutions

OpenStudy (anonymous):

Ohh, okay thanks again

jimthompson5910 (jim_thompson5910):

np